OpenStudy (anonymous):

Find a basis for the subspace S of all diagonal 3  3 matrices of R^3x3

OpenStudy (jamesj):

Well, in general a diagonal matrix looks like this: \[\left[\begin{matrix}a & 0 & 0 \\ 0 & b & 0 \\ 0 & 0 & c \end{matrix}\right]\]

OpenStudy (jamesj):

So hopefully it's not clear what three matrices would form a natural basis \[a \times vector1 + b \times vector2 + c \times vector3\]